OpenStudy (jared8651):

The half life of a Rock is: 6.7 Billion years Amount Element: 10% Amount lead: 90% Age of Rock: ? Can someone please help?

ganeshie8 (ganeshie8):

90% lead tells us that 90% of the rock has been decayed so far

OpenStudy (jared8651):

ok.

ganeshie8 (ganeshie8):

Half life of 6.7 billion years tells us that half of the stuff gets decayed every 6.7 billion years : \[Q' = Q*\left(\dfrac{1}{2}\right)^{t/6.7}\] Observe that when t = 6.7 in above equation, you do get Q' = Q/2 as expected

OpenStudy (jared8651):

What does Q represent?

ganeshie8 (ganeshie8):

Q is the starting `amount of substance` Q' is the substance remaining after `t` billion years

ganeshie8 (ganeshie8):

Simply replace Q' by 0.1Q and solve t

ganeshie8 (ganeshie8):

\[0.1Q = Q*\left(\dfrac{1}{2}\right)^{t/6.7}\]

OpenStudy (jared8651):

Can you please show me the process im still confused.

ganeshie8 (ganeshie8):

Start by cancelling Q both sides

OpenStudy (jared8651):

ok

ganeshie8 (ganeshie8):

\[0.1 = \left(\dfrac{1}{2}\right)^{t/6.7}\] Your goal is to isolate \(t\)

ganeshie8 (ganeshie8):

How good are you with logarithms ?

OpenStudy (jared8651):

not good at all

OpenStudy (jared8651):

Thank you so much for your help

ganeshie8 (ganeshie8):

This is the best example to make sense of logarithms too

ganeshie8 (ganeshie8):

Let me introduce you to logarithms quick. Look at the expression \(2^3 = 8\). Is it a true statement ?

OpenStudy (jared8651):

hmm

OpenStudy (jared8651):

yes

ganeshie8 (ganeshie8):

Focus on the exponent. The thing on the top : 3.

ganeshie8 (ganeshie8):

\(2^3 = 8\) We say \(3\) is the logarithm of 8 when base is 2.

ganeshie8 (ganeshie8):

As you can see, logarithm = exponent

ganeshie8 (ganeshie8):

Look at \(10^4 = 1000\) What is the logarithm of 1000 in base 10 ?

OpenStudy (jared8651):

oh ok its starting to make scense

OpenStudy (jared8651):

\[10^3?\]

ganeshie8 (ganeshie8):

Careful, logarithm = exponent

ganeshie8 (ganeshie8):

Only exponent

OpenStudy (astrophysics):

 Expressed mathematically, this might be a bit helpful